OpenStudy (anonymous):

The solution to the following equation is 2.75 2x - 2 / 5 = 0.75 TRUE OR FALSE?

OpenStudy (anonymous):

f

OpenStudy (anonymous):

2x-2/5 =0.75 2x = 0.75 + 2/5 2x = 0.75 + 0.4 x=1.15/2 = 0.575 2.75 is False